OpenStudy (anonymous):

A volume of 18.0L contains a mixture of 0.250mole N2, 0.250mole O2, and an unknown quantity of He. The temperature of the mixture is 0 degree C, and the total pressure is 1.00atm. How many grams of helium are present in the gas mixture?

OpenStudy (mos1635):

P(total) * V = n(total)*R*T n(total)=n1+n2+ n3 n3=m3/MM3